OpenStudy (anonymous):

Medals & fan anybody ? Given the following segment, find the value of z. a. 8 b. –7 c. 7.5 d. –8

OpenStudy (mokeira):

Which segment

OpenStudy (mokeira):

\[(4z-15)=(z-3)+(z+4)\] \[4z-15= 2z+1\] \[4z-2z=1+\] \[2z=16\] \[z=16\] z=8

OpenStudy (anonymous):

thankyou :)